Bob invested some money at 6% and $300 less at 8%. His total annual interest income is $116. How much did he invest at 6%?
Answer by nerdybill(7384)

Bob invested some money at 6% and $300 less at 8%. His total annual interest income is $116. How much did he invest at 6%?
.
Let x = amount ($) invested at 6%
then
x-300 = amount invested at 8%
.
.06x + .08(x-300) = 116
.06x + .08x-24 = 116
.14x-24 = 116
.14x = 140
x = $1000